OpenStudy (anonymous):

if f(x) is a linear function such that f(1)=5 and f(4)=-7,then what is the value if f(-1)?

OpenStudy (anonymous):

you know two points (1,5) (4,7) find slope and line equation

OpenStudy (anonymous):

then evaluate at -1

OpenStudy (anonymous):

imranmeah91 may I please have help.

OpenStudy (anonymous):

Knowing 2 Y coordinates and 2 X coordinates you can get the slope of the line using \[M = (y_{2} - y_{1}) \div (x_{2} - x_{1})\] M = 2/3 Now knowing that we can solve for the y intercept 5 = (2/3) * 1 + b Solve for b B = 13/3 Now we know b and m, plug that into the equation for a line y=mx +b y = ((2/3)*-1) + 13/3 y = (-2/3) +13/3 y = 11/3 Therefore f(-1) = 11/3